# Sharing Matters: Analysing Neurons Across Languages and Tasks in LLMs

### v2

@@ 6.3 The Effectiveness of Neuron Types
| w/o. all-shared | 9.85% | 16.7 | 3.5 | 10.1 | 10.0 | 6.6 | 9.0 | 1.4 | 12.1 | 14.5 |
| w/o. partial-shared | 10.30% | 52.9 | 40.4 | 49.7 | 47.6 | 49.2 | 40.3 | 36.1 | 50.0 | 50.0 |
| w/o. specific | 3.12% | 53.7 | 41.7 | 50.3 | 48.9 | 47.4 | 40.6 | 35.3 | 50.4 | 49.3 |
| w/o. non-activated | 76.73% | 36.6 | 31.6 | 33.6 | 33.4 | 29.5 | 31.3 | 28.3 | 34.5 | 23.5 |
@@
| baseline | 0% | 96.2 | 55.1 | 48.9 | 49.5 | 30.6 | 6.3 | 38.8 | 51.7 | 33.4 |
| w/o. all-shared | 8.69% | 11.0 | 4.9 | 6.1 | 4.9 | 1.9 | 0.4 | 1.5 | 6.1 | 2.9 |
| w/o. partial-shared | 13.20% | 90.2 | 51.7 | 48.9 | 48.9 | 25.4 | 5.5 | 35.5 | 50.9 | 38.4 |
| w/o. specific | 4.87% | 96.1 | 54.4 | 48.7 | 48.9 | 30.4 | 6.3 | 37.9 | 51.7 | 28.5 |
| w/o. non-activated | 73.20% | 36.1 | 15.8 | 18.2 | 17.8 | 1.9 | 9.8 | 19.9 | 10.6 | 16.3 |
@@
| w/o. random | 15% | 94.8 | 47.1 | 36.1 | 39.4 | 0.8 | 4.3 | 28.4 | 40.4 | 11.1 |
| w/o. random | 25% | 91.5 | 46.8 | 36.2 | 38.6 | 1.1 | 4.4 | 27.9 | 40.4 | 12.1 |
| w/o. random | 75% | 11.1 | 5.5 | 9.3 | 11.3 | 0.1 | 2.7 | 1.9 | 8.9 | 7.1 |
7 Verification on Downstream Tasks
We analyse the behaviors of different neuron types in Section 5 and explore the contribution and effectiveness of various types of neurons in Section 6, demonstrating that neuron shared across languages and tasks are of vital importance from an internal representation perspective. In this section, we conduct experiments on downstream tasks to verify our analytical findings regarding the significance of neuron sharing.
We assess accuracy changes by selectively deactivating four specific types of neurons along with a control group of randomly selected neurons, setting their activation values to zero. These experiments are performed using the XNLI and Cross-lingual KE (EN (Edit) ALL (Test)) tasks on the BLOOMZ-7b model, with results detailed in Table 2. The most significant decrease in accuracy is observed when the all-shared neurons are deactivated (e.g. 91.6% decrease in the XNLI German testset). While the deactivation of specific and partial-shared neurons also impacts accuracy, these effects are less pronounced compared to those of all-shared neurons. Deactivating neurons randomly has minimal impact on the performance. In order to prove the key role of all-shared neurons across different LLMs, we conduct the ablation experiments with LLAMA2-7b-chat backbone in Appendix A.6.1.
@@ 8 Conclusion
In this study, we explored the complex mechanisms of neuron activation within multilingual LLMs, addressing the significant research gap in understanding these models beyond a monolingual context. We developed a fine-grained classification for analysing how neurons respond to different tasks and languages. We categorized neurons into four distinct groups: all-shared, partial-shared, specific, and non-activated. Our research revealed that neuron sharing is heavily influenced by the specific characteristics of tasks and examples, and, surprisingly, it does not always align with language similarity. Furthermore, neurons shared across all languages proved essential for generating accurate responses, highlighting their pivotal role in multilingual processing. Our study improves the understanding of the internal workings of multilingual LLMs and fosters future research in this direction.
9 Limitations
In this paper, we develop a method to analyse neuron behaviors in detail by categorizing them into four distinct neuron types w.r.t the degree of their responses to input languages. Although this enables a fine granularity neuron analysis on LLM backbones across various linguistic characteristics and task complexity, the scope of the experiments can be extended to accommodate larger LLMs with large amounts of parameters (i.e., BLOOMZ-176b) on a more comprehensive range of tasks. While this study demonstrates that the number of languages slightly impacts the percentage of all-shared neurons, it is limited to 10 languages. Exploring the effects of incorporating a larger number of languages into the proposed method warrants further investigation. Moreover, while all-shared neurons in FFN blocks are identified to be of great importance, the specific mechanisms by which they operate remain unclear. Additionally, other network components, for example, attention heads, are not in the scope of this analysis.
Appendix A Appendix
A.1 Detailed Interpretation of Projection in Vocabulary Space
@@
A.2 Tasks
XNLI. Natural Language Inference (Conneau et al., 2018) is a multilingual natural languages inference dataset, containing 5000 items. Each test sample consists of a premise and a hypothesis, requiring an LLM to determine whether a hypothesis is entailed, contradicted, or neutral conditioned on the premise.
Fact Probing. LLMs are used to predict factual answers in response to corresponding probing prompts. A multilingual factual knowledge dataset (mParaRel (Fierro and Søgaard, 2022)) capturing 38 binary relations (e.g., X born-in Y) is used in the analysis. We seletc the relation of “capital” subset (X capital Y) as testset, including 348 items.
Cross-lingual Knowledge Editing (KE). MzsRE (Wang et al., 2023) is a multilingual question-answering dataset, containing 743 items for each language. It provides counterfactual edited knowledge in the context and requires an LLM to produce the corresponding answer according to the context. We evaluate LLMs in two Cross-lingual KE scenarios: 1) EN (Edit) ALL (Test): edit in English and test in other languages and 2) ALL (Edit) EN (Test): edit in other languages and test in English.
A.3 Prompts

### v3

@@ 5 Shared Neurons Are Crucial to Performance
| w/o. all | 9.9% | 16.7 | 3.5 | 10.1 | 10.0 | 6.6 | 9.0 | 1.4 | 12.1 | 14.5 |
| w/o. partial | 10.3% | 52.9 | 40.4 | 49.7 | 47.6 | 49.2 | 40.3 | 36.1 | 50.0 | 50.0 |
| w/o. specific | 3.1% | 53.7 | 41.7 | 50.3 | 48.9 | 47.4 | 40.6 | 35.3 | 50.4 | 49.3 |
| w/o. non-act. | 76.7% | 36.6 | 31.6 | 33.6 | 33.4 | 29.5 | 31.3 | 28.3 | 34.5 | 23.5 |
@@
| baseline | 0% | 96.2 | 48.8 | 36.9 | 49.5 | 24.6 | 6.3 | 38.8 | 49.4 | 33.4 |
| w/o. all | 8.7% | 11.0 | 4.9 | 6.1 | 4.9 | 1.9 | 0.4 | 1.5 | 6.1 | 2.9 |
| w/o. partial | 13.3% | 90.2 | 51.7 | 46.9 | 48.9 | 25.4 | 5.5 | 35.5 | 50.9 | 38.4 |
| w/o. specific | 4.9% | 96.1 | 54.4 | 48.7 | 48.9 | 30.4 | 6.3 | 37.9 | 51.7 | 28.5 |
| w/o. non-act. | 73.2% | 36.1 | 15.8 | 18.2 | 17.8 | 1.9 | 9.8 | 19.9 | 10.6 | 16.3 |
@@
| 15% | 94.8 | 47.1 | 36.1 | 39.4 | 0.8 | 4.3 | 28.4 | 40.4 | 11.1 | |
| 25% | 91.5 | 46.8 | 36.2 | 38.6 | 1.1 | 4.4 | 27.9 | 40.4 | 12.1 | |
| 75% | 11.1 | 5.5 | 9.3 | 11.3 | 0.1 | 2.7 | 1.9 | 8.9 | 7.1 | |
| Cross-lingual KE (ALL (Edit) EN (Test)) task | ||||||||||
| baseline | 0% | 96.2 | 55.1 | 49.2 | 49.5 | 30.6 | 9.2 | 39.3 | 51.7 | 36.6 |
| w/o. all | 10.2% | 24.4 | 19.5 | 13.8 | 13.1 | 8.0 | 1.4 | 14.5 | 19.9 | 7.1 |
| w/o. partial | 10.5% | 85.1 | 51.3 | 47.8 | 48.0 | 25.4 | 5.2 | 35.1 | 51.4 | 36.1 |
| w/o. specific | 3.8% | 96.1 | 54.4 | 48.7 | 48.9 | 29.7 | 6.3 | 38.0 | 51.8 | 30.0 |
| w/o. non-act. | 75.5% | 19.2 | 19.8 | 15.5 | 14.7 | 11.4 | 2.0 | 18.2 | 12.0 | 7.5 |
| w/o. random | 5% | 95.6 | 54.2 | 48.7 | 50.2 | 29.9 | 6.5 | 38.5 | 51.3 | 33.0 |
| 15% | 93.9 | 56.1 | 49.1 | 49.9 | 29.2 | 6.4 | 38.2 | 51.0 | 32.6 | |
| 25% | 91.3 | 55.9 | 48.5 | 49.3 | 28.3 | 6.8 | 37.7 | 50.3 | 29.7 | |
| 75% | 10.2 | 17.0 | 11.7 | 15.7 | 4.7 | 1.1 | 7.8 | 14.8 | 7.0 | |
| Fact Probing task | ||||||||||
| baseline | 0% | 72.4 | 41.6 | 56.6 | 58.1 | 37.3 | 5.7 | 39.3 | 57.4 | 51.4 |
| w/o. all | 0.2% | 43.4 | 12.9 | 34.4 | 22.4 | 11.4 | 5.2 | 15.2 | 34.5 | 29.0 |
| w/o. partial | 36.7% | 43.3 | 20.8 | 31.2 | 30.9 | 14.4 | 2.8 | 24.5 | 34.6 | 29.4 |
| w/o. specific | 16.6% | 18.1 | 9.1 | 30.1 | 7.7 | 5.7 | 5.7 | 9.4 | 12.3 | 22.1 |
| w/o. non-act. | 46.4% | 42.5 | 27.6 | 39.9 | 28.1 | 1.7 | 0.0 | 22.9 | 40.2 | 17.5 |
| w/o. random | 1% | 76.4 | 50.6 | 48.6 | 56.0 | 3.2 | 0.0 | 36.2 | 59.5 | 47.1 |
| 15% | 71.5 | 48.5 | 45.6 | 63.5 | 4.3 | 0.0 | 22.7 | 44.5 | 38.2 | |
| 35% | 77.3 | 51.4 | 50.3 | 56.3 | 4.6 | 0.0 | 37.1 | 57.5 | 48.3 | |
| 45% | 29.0 | 21.3 | 16.4 | 17.2 | 0.3 | 0.0 | 9.5 | 25.9 | 6.0 | |
In this section, we explore how different neuron types affect the performance of the BLOOMZ-7b model by selectively deactivating specific groups of neurons. By setting the activation values of these neurons to zero, we assess their impact on the model’s output across various tasks. Specifically, we compare the effects of deactivating four distinct types of neurons and include a control group of randomly selected neurons to evaluate their respective contributions to the model performance. Our experiments involve tasks such as XNLI, cross-lingual KE, and fact probing.
All-shared neurons play a crucial role in model performance across different tasks.
As shown in Table 1, we observe that all-shared neurons significantly contribute to the model’s performance across various tasks. For instance, in the Cross-lingual KE (EN (Edit) ALL (Test)) task, deactivating the all-shared neurons, which account for only 8.71% of the total neurons, results in an 87.39% decrease in accuracy. Moreover, for the Fact Probing task, deactivating the all-shared neurons, which constitute only 0.28% of the total neurons, causes a substantial 50.31% performance drop. Furthermore, deactivating the specific neurons, which account for 16.56% of the total neurons, leads to the largest performance decline of 67.41%. In comparison, deactivating a comparable number of random selected neurons typically results in smaller performance drops, suggesting that all-shared neurons are crucial to the performance.
Deactivating neurons does not always result in performance declines.
Interestingly, we sometimes observe small performance gains when a small number of neurons are deactivated, as shown in Table 1, regardless of the neuron type. To explore this phenomenon further, we provide a breakdown of the results by language in Table 2. Our analysis reveals that only deactivating the all-shared neurons consistently leads to a decline in model performance across various tasks and languages. In contrast, deactivating either partial-shared neurons or specific neurons can occasionally improve performance for certain languages. For example, in the Cross-lingual KE (EN (Edit) ALL (Test)) task, we observe substantial performance improvements in German (de), Spanish (es), and Chinese (zh) when the partial-shared neurons are deactivated. We hypothesize that this phenomenon stems from knowledge conflicts encoded in the LLM (Xu et al., 2024). By deactivating certain neurons, these knowledge conflicts may be mitigated, resulting in enhanced performance.
It is important to note that we conduct similar experiments using the LLAMA2-7b-chat and present the results in Appendix E. These additional experiments yield observations and conclusions consistent with those using BLOOMZ-7b.
6 Probing Neuron Contributions
We demonstrate the significant role of neurons shared across languages, particularly the all-shared neurons, in generating answers, as discussed in Section 5. To gain a deeper understanding of the model’s behavior, we conduct further analysis using two metrics: Generation Impact Score and Correctness Impact Score. First, we introduce the definitions of these two metrics in Section 6.1. Then, we analyse and quantify the contributions of each type of neuron in Section 6.2 and Section 6.3, respectively.
6.1 Generation Impact Score and Correctness Impact Score
In this section, we introduce two measures to quantify the contributions of neurons: Generation Impact Score and Correctness Impact Score.
Generation Impact Score
Inspired by Geva et al. (2022), the Generation Impact Score () evaluates the importance of neurons in generating answers. For the -th neuron at -th layer, the is defined as:
which is the proportion of its weight to the sum of weights of all neurons in the FFN block. is the absolute value of activation value and is the L2-norm of value (the column in matrix ).
Correctness Impact Score
Following Geva et al. (2022) and Voita et al. (2023), Correctness Impact Score () assesses a neuron’s influence on generating the correct answer.
where is the embedding of the correct answer . A larger has a higher probability to produce the correct answer , while a negative reduces the probability in generating . Detailed descriptions of the neuron projection are provided in Appendix B.
| all-shared | partial-shared | specific | non-activated | |||||||||||||
| max | min | mean | var | max | min | mean | var | max | min | mean | var | max | min | mean | var | |
| en | 1.85 | -0.94 | 0.07 | 0.36 | 0.22 | -0.16 | 1.2e-4 | 1.9e-4 | 0.02 | -0.02 | 2.5e-4 | 3.5e-5 | 0.04 | -0.03 | 2.1e-4 | 5.8e-6 |
| de | 1.03 | -0.60 | 0.02 | 0.07 | 0.13 | -0.13 | 6.7e-5 | 7.1e-5 | 0.07 | -0.03 | 2.3e-5 | 2.3e-5 | 0.02 | -0.01 | 2.9e-6 | 2.9e-6 |
| es | 1.15 | -0.84 | 0.02 | 0.07 | 0.12 | -0.11 | 1.3e-4 | 6.3e-5 | 0.01 | -0.01 | 9.4e-5 | 7.6e-6 | 0.02 | -0.02 | 7.7e-5 | 3.1e-6 |
| fr | 1.06 | -0.78 | 0.01 | 0.05 | 0.15 | -0.11 | 2.1e-4 | 7.8e-5 | 0.03 | -0.04 | 3.6e-5 | 9.9e-6 | 0.02 | -0.02 | 5.6e-5 | 2.8e-6 |
| ru | 0.70 | -0.45 | 3.3e-3 | 5.9e-3 | 0.24 | -0.13 | 2.6e-4 | 7.6e-5 | 0.08 | -0.03 | 1.1e-4 | 1.8e-5 | 0.01 | -0.01 | 3.9e-5 | 1.7e-6 |
| th | 0.50 | -0.90 | 2.6e-3 | 0.02 | 0.17 | -0.10 | 2.2e-4 | 6.8e-5 | 0.03 | -0.05 | 3.4e-5 | 1.8e-5 | 0.01 | -0.01 | 7.1e-5 | 1.9e-6 |
| tr | 0.82 | -0.51 | 0.03 | 0.07 | 0.12 | -0.12 | 1.6e-4 | 7.4e-5 | 0.04 | -0.03 | 6.6e-5 | 1.1e-5 | 0.02 | -0.02 | 9.1e-5 | 3.4e-6 |
| vi | 0.86 | -0.68 | 6.8e-3 | 0.03 | 0.15 | -0.11 | 9.3e-5 | 6.8e-5 | 0.04 | -0.04 | 2.8e-5 | 1.3e-5 | 0.02 | -0.02 | 3.2e-5 | 2.6e-6 |
| zh | 0.52 | -0.42 | 1.9e-3 | 0.02 | 0.17 | -0.20 | 1.5e-4 | 7.7e-5 | 0.08 | -0.07 | 8.5e-5 | 2.6e-5 | 0.02 | -0.01 | 1.7e-5 | 3.1e-6 |
Comparison
While both Generation Impact Score () and Correctness Impact Score () measure neuronal influence, they serve different purposes. The quantifies a neuron’s overall contribution to the generation process, regardless of output correctness. In contrast specifically measures a neuron’s impact on producing accurate responses by incorporating the correct answer’s embedding. Thus, the key distinction lies in their consideration of answer correctness: focuses on general generation ability, whereas emphasizes correctness.
6.2 The Generation Impact of Neuron Types
In this section, we explore the contribution of each neuron type using the Generation Impact Score () described in Section 6.1.
All-shared neurons have the greatest impact on generation outputs.
As shown in Figure 3(a), we analyse the across layers on the English and German test sets of three tasks (with overall results provided in Appendix F). For both English and German, it can be observed that the all-shared neurons almost always achieve the highest across all layers, indicating their significant influence on the model’s output generation. The partial-shared neurons are the second most influential, particularly in the upper layers. Notably, there is a decrease in the influence of all-shared neurons between layers 5 and 10. This can be attributed to the fact that assesses the impact on generating answers, while the lower layers are primarily responsible for input understanding (Zhao et al., 2024b). Consequently, all types of neurons exhibit lower in these layers. Moreover, previous studies have demonstrated that higher layers capture more abstract, high-level information essential for generation (Gao et al., 2024). These findings suggest that shared neurons play a more significant role in the model’s generation capabilities.
6.3 The Correctness Impact of Neuron Types
In this section, we assess the effectiveness of each neuron type using the Correctness Impact Score () described in Section 6.1.
All-shared neurons have the greatest impact on generating correct answers.
In the Cross-lingual KE (EN (Edit) ALL (Test)) task, we present the maximum, minimum, average, and variance of for each neuron type across all layers of the BLOOMZ-7b model, as shown in Table 3. The results reveal that all-shared neurons have both the highest maximum and the lowest minimum values, indicating that they have strong impact on generating correct outputs. While all-shared and partial-shared neurons display a wide variance in (e.g., 1.85 vs. -0.94 and 0.22 vs. -0.16 in English, respectively), specific neurons and non-activated neurons exhibit much narrower score ranges (approximately 0.07). Furthermore, the all-shared neurons also exhibit the largest mean and variance of among all kinds of neurons.
In conclusion, these findings presented in Section 6.2 and Section 6.3 demonstrate that all-shared neurons also have the greatest impact on generating both answers and correct answers, highlighting their importance in the model’s performance across different languages and tasks.
7 Understanding Neuron Activations
We demonstrate in Section 5 that shared neurons have a significant impact on model performance and investigate their influence on the generation process in Section 6. However, the inner patterns of neurons across layers remain unexplored. In this section, we firstly introduce the measure of quantifying neuron activation in Section 7.1, and then we further illustrate how neuron activation patterns vary across tasks (Section 7.2), LLMs (Section 7.3) and languages (Section 7.4).
7.1 Measuring Neuron Activation
In this section, we explain how to quantify neuron activation patterns based on the definitions in Section 3.2. Specifically, we measure the percentage of each type of neuron relative to the total number of neurons. Given the -th test instance, the percentage of each neuron type at the -th layer can be defined as follows:
where denotes the number of elements in the set. Consequently, the aggregated neuron activation pattern at the -th layer for one dataset containing instances can be defined as:
7.2 Neuron Activations Across Tasks
Neuron activations are task-related.
As shown in Figure 3, non-activated neurons are typically more prevalent than other types of neurons, except in the Fact Probing task. In this task, there are more partial-shared neurons and specific neurons, with a negligible amount of all-shared neurons, whereas other tasks involve far more all-shared neurons. Referring to Table 1, deactivating the specific neurons and all-shared neurons results in the largest and second largest performance declines. These findings demonstrate that the some factual knowledges in LLMs are language-specific and minimally shared across languages, while others are universally shared. We leave more in-depth investigation to the future work.
Neuron sharing peaks at early layers for universal features, declining later for specific ones.
We present the percentage of each neuron type at each layer in Figure 3. The number of all-shared neurons and partial-shared neurons typically peaks between the 5th and 10th layers and then gradually decreases in subsequent layers. This trend can be explained by the functional roles of different layers in the model. The initial layers, which are closer to the input data, primarily focus on capturing low-level features such as basic lexical and syntactic patterns. As the network progresses to the later layers (between the 5th and 10th layers), it begins to learn abstract concepts that are relatively universal across different tasks and languages. This universality leads to a higher number of shared neurons in these layers. In contrast, the higher layers specialize in task-specific features and nuances unique to each task, resulting in a decline in neuron sharing. These findings highlight the importance of neuron sharing in LLMs, as shared neurons in the early layers facilitate the transfer of universal knowledge across tasks and languages. They also align with previous research (Yosinski et al., 2014; de Vries et al., 2020; Zhao et al., 2024b; Gao et al., 2024).
7.3 Neuron Activations Across LLMs
Different LLMs exhibit different neuron activation patterns.
To investigate whether neuron activation patterns vary across different multilingual LLMs, we present additional results from LLaMA2-7b-chat in Figure 4. Our analysis reveals that the activation patterns in LLaMA2-7b-chat differ significantly from those observed in BLOOMZ-7b, highlighting the variability across models. Notably, LLaMA2-7b-chat demonstrates a higher degree of neuron sharing, particularly for partial-shared neurons. This phenomenon can be attributed to the English-centric nature of LLaMA2-7b-chat. When processing multilingual inputs, the model heavily relies on knowledge transfer from English to other languages, resulting in a substantial number of partial-shared neurons. We also present additional results using XGLM (Lin et al., 2022) in Appendix G, aligning with our observations.
Instruction finetuned LLMs exhibit larger proportion of the all-shared neurons.
We conduct additional experiments using the foundation model BLOOM-7b to explore the impact of instruction finetuning on neuron activation patterns. As shown in Figure 5, the instruction-finetuned BLOOMZ-7b demonstrates a higher percentage of all-shared neurons compared to BLOOM-7b. This observation suggests that instruction finetuning may encourage neuron sharing within LLMs, potentially aligning their internal representations across languages. Therefore, instruction-finetuned LLMs, such as BLOOMZ-7b, generally outperform their foundational counterparts.
7.4 Neuron Activations Across Languages
Neuron sharing does not completely align with language similarity.
We investigate the relationship between language similarity and neuron sharing by analysing the proportion of partial-shared neurons for language pairs involving German and several other languages on the Fact Probing task. As shown in Figure 6, our findings reveal that similar languages (e.g., German and French) do not always exhibit higher levels of neuron sharing. For instance, the proportion of partial-shared neurons between German and Chinese is nearly identical to that between German and French, despite German and French both belonging to the Indo-European language family, while Chinese belongs to the Sino-Tibetan language family. Furthermore, we observe no consistent pattern in the percentage of specific neurons across the languages studied, suggesting that neuron specialization may not directly correlate with language similarity. We leave further exploration of this phenomenon to future work. Additional results on XNLI task are in Appendix I.
Furthermore, we conduct ablation studies to investigate the impact of two key factors on the neuron activation patterns: the size of the backbone model with 0.56b, 1b, 3b, 7b parameters (Appendix J.).
@@ 8 Conclusion
In this study, we explored the complex mechanisms of neuron activation within multilingual LLMs, addressing the significant research gap in understanding these models beyond a monolingual context. We developed a fine-grained classification for analysing how neurons respond to different tasks and languages. We categorized neurons into four distinct groups: all-shared, partial-shared, specific, and non-activated. Our research revealed that neurons shared across all languages proved essential for generating accurate responses, highlighting their pivotal role in multilingual processing. Furthermore, we demonstrate that neuron sharing is task-related, and, it does not always align with language similarity. Our study improves the understanding of the internal workings of multilingual LLMs and fosters future research in this direction.
Acknowledgement
This work is funded by EU Horizon Europe (HE) Research and Innovation programme grant No 101070631, and UK Research and Innovation under the UK HE funding grant No 10039436.
The computations described in this research were performed using the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk/). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
Appendix A Detailed Interpretation of Projection in Vocabulary Space
@@
Appendix B Tasks
XNLI. Natural Language Inference is a multilingual natural languages inference dataset, containing 5000 items. Each test sample consists of a premise and a hypothesis, requiring an LLM to determine whether a hypothesis is entailed, contradicted, or neutral conditioned on the premise.
Fact Probing. LLMs are used to predict factual answers in response to corresponding probing prompts. A multilingual factual knowledge dataset (mParaRel) capturing 38 binary relations (e.g., X born-in Y) is used in the analysis. We seletc the relation of “capital” subset (X capital Y) as testset, including 348 items.
Cross-lingual Knowledge Editing (KE). MzsREis a multilingual question-answering dataset, containing 743 items for each language. It provides co unterfactual edited knowledge in the context and requires an LLM to produce the corresponding answer according to the context. We evaluate LLMs in two Cross-lingual KE scenarios: 1) EN (Edit) ALL (Test): edit in English and test in other languages and 2) ALL (Edit) EN (Test): edit in other languages and test in English.
Appendix C Prompts